HIGH RESOLUTION ANION THRESHOLD PHOTODETACHMENT SPECTROSCOPY BY VELOCITY MAP IMAGING OF SLOW PHOTOELECTRONS

\underline{ANDREAS OSTERWALDER}, MATHEW J. NEE, JIA ZHOU, AND DANIEL M. NEUMARK, Department of Chemistry, University of California, Berkeley, CA 94720 (USA).

We have combined Velocity Map Imaging (VMI) and threshold photodetachment (TPD) to obtain a new method for high resolution anion photoelectron spectroscopy. VMI is used to detect slow photoelectrons, generated by photodetachment in a tunable laser. With a maximum kinetic energy of the detected photoelectrons of ca. 10-50 meV and an energy resolution of better than ΔE/E=5 %, we achieve an ultimate resolution of a few cm^{−1}. The sensitivity of the technique is guaranteed by the collection efficiency of VMI, which is close to 100 %, and by the possibility of detecting photoelectrons with low but non-zero kinetic energy, whereby limitations by the Wigner threshold law can be avoided. This renders VMI-TPD spectroscopy more generally applicable than conventional anion zero kinetic energy photoelectron spectroscopy while maintaining the high resolution. We have applied the new technique to the investigation of Cl-D_{2}. Franck-Condon overlap between the wave functions of anion and neutral allows the investigation of the shallow potential well located near the transition state of the Cl+H_{2} benchmark chemical reaction. This well is deep enough to contain a low number of van der Waals bound states and has been observed in previous photoelectron studies carried out in our lab. At a resolution of ca. 8 meV, these spectra revealed an increased linewidth of Cl-D_{2} vs. Cl-H_{2}. This finding was explained by hindered rotor structure that is observable only in the deuterium cluster but could not be fully resolved by conventional PES. The increased resolution of VMI-TPD allows the resolution of this, and potentially even finer, structure and will help toward a complete understanding of the dynamics of this reaction.